1. If the demand for a particular brand of car decreases by 2% due to an increase in income by 10%, how much is the income elasticity? Is this product normal/inferior good?

2. Suppose that when the average family income rises from $30,000 per year to $40,000 per year, the average family’s purchases of toilet paper rise from 100 rolls to 105 rolls per year. Use mid-point formula. (circle the correct answers).

  1. The income-elasticity of demand for toilet paper is -0.17/      0.17/     -0.27/   0.27.
  2. The demand for toilet paper is income-inelastic/elastic.
  3. Toilet paper is a normal necessity/normal luxury/inferior.
Answer #1

a) If the demand decrease by 2% and the income increase is 10%. income elasticity is -2/10 = 0.2 and the good is a normal good.

b) Income elasticity = %change in demand / %change in income

% change in demand = (105 -100 ) / ((100+105) / 2)

= 5 / 102.5

= 0.04

% change in income = (40,000 -30,000) / ((40,000+30,000)/2)

10,000 / 35,000

= 0.28.

Income elasticity = 0.04 / 0.28 = 0.145 closest to 0.17. The answer is "B". The demand for toilet paper is inelastic and it is a normal good.